(57+7)+(2 +3 +4 k, where t o Find T1-1). (20) Consider the curver() NI-1) and B(-1)
Answer #1

Given Curve, 8+) = (57+7)i + (21 +3)j +(4-Ė)K -OctZoo x(A) = [(54+7) i +(2813)j +(4-2)x7 r(t) = 5 i +4t j - 2 tk (A) = f ( 2TC+ (4) 18() si + 4 + ;-2&K 5° +(4+)? + (-24) 5i + 4 + ;-24k - 25 + 16 +² +42 si +4t; -24k √ 25+ 204 = T(A) = 47 TA +2012 25+How T(4) 5i +4 tj - 2 tk ) 125+2012 Vest2012 Sep (51+4+3 =24) – (si+4+=241Jesus J25+20fe)? 20 125+20 + (01 +432k) -(si+4+j=#)(0-1004)i + (100+804-80) ; +(-50-70 tok (25+2042 / T + -loot i +10oj - 50k (25 +20 17% ht : T(H) = toox(-1) i tlooj 50 k (25+T(-1) NH = TT- ((-/) [T(-)) (95* trooi +100j-50K) 150 (4522 looi tlooj - 50k 150 M(-) = i + k and B(-1) = T(-1) X N(-1)() 0 + 1+、 a se) = i + t = j+后 i +=k 12 (BA) =